# LoLA: Low-Rank Linear Attention With Sparse Caching

### v1

@@ Abstract
Transformer-based large language models suffer from quadratic complexity at inference on long sequences. Linear attention methods are efficient alternatives, however, they fail to provide an accurate approximation of softmax attention. By additionally incorporating sliding window attention into each linear attention head, this gap can be closed for short context-length tasks. Unfortunately, these approaches cannot recall important information from long contexts due to “memory collisions”. In this paper , we propose LoLA: Low-rank Linear Attention with sparse caching. LoLA separately stores additional key-value pairs that would otherwise interfere with past associative memories. Moreover, LoLA further closes the gap between linear attention models and transformers by distributing past key-value pairs into three forms of memory: (i) recent pairs in a local sliding window; (ii) difficult-to-memorize pairs in a sparse, global cache; and (iii) generic pairs in the recurrent hidden state of linear attention. As an inference-only strategy, LoLA enables pass-key retrieval on up to 8K context lengths on needle-in-a-haystack tasks from RULER. It boosts the accuracy of the base subquadratic model from 0.6% to 97.4% at 4K context lengths, with a 4.6$\times$ smaller cache than that of Llama-3.1 8B. LoLA demonstrates strong performance on zero-shot commonsense reasoning tasks among 1B and 8B parameter subquadratic models. Finally, LoLA is an extremely lightweight approach: Nearly all of our results can be reproduced on a single consumer GPU.
@@ 1 Introduction
Transformer-based large language models (LLMs) rely on storing all past tokens in an ever-growing key-value (KV) cache [47]. This allows future query tokens to access past memories with associative recall, which enables in-context learning [32]. Since no previous information is discarded, the KV cache continues to grow with context length. This eventually leads to a memory bottleneck on long context tasks. As a result, transformers cannot condition next token predictions on arbitrarily long sequences.
Alternative architectures to transformers have been proposed—such as Mamba [21], DeltaNet [41], linear attention [27], and others [51, 3, 45]—to reduce the compute complexity from quadratic to linear. Additionally, these approaches reduce the memory cost from linear to constant. In particular, linear attention removes the exponential dot product in softmax [27]. This effectively collapses the unbounded KV-cache into a fixed-size matrix, which corresponds to a recurrently formed hidden state (i.e., a linear RNN). This constructs a linear associative memory map from keys to values. Past memories can be recalled through a vector-matrix product of an incoming query vector and the hidden state matrix. Linear attention enables constant-cost prediction per token when conditioned on arbitrarily long contexts.
While efficient and flexible, linear attention architectures lag behind transformers in language modeling tasks [52]. This is largely noticeable on tasks leveraging in-context learning [33, 25]. The removal of the exponential dot product allows for non-orthogonal keys to interfere with the hidden state’s learned key-to-value map. This interference impairs associative recall. Previous work used nonlinear query and key activations to improve the exponential dot product approximation [11, 57]. Though, these attempts are essentially performing a low-rank approximation of the infinite-rank exponential dot product kernel.
Additional use of sparse attention [10] can improve linear attention’s low-rank approximation. Previously, LoLCATs [56] exploited locality in natural language by augmenting feature-map-based linear attention with sliding-window attention [1]. Furthermore, this method demonstrates that pretrained Transformers can be distilled into linear attention architectures (see also the concurrent work [49, 6, 5]). The use of distillation and sliding-window attention performs well for short-context tasks. These models, however, struggle to recall critical, long-term facts that fall outside the sliding window.
@@
How can associative memory for language models be improved with a finite memory footprint?
@@
We present LoLA: Low-rank Linear Attention with sparse caching. LoLA distributes historical tokens into three forms of memory: (i) recent KV pairs are stored in a sliding window cache, (ii) difficult-to-memorize pairs in a sparse global cache, and (iii) all other pairs are placed in a recurrent hidden-state matrix via linear attention. LoLA performs a self-recall check to see which KV pairs disagree with the current hidden state’s linear associative map. LoLA sparsely caches the interfering memories in full-rank. The selection mechanism effectively mitigates memory collisions with a small, constant-sized cache. This inference strategy can be applied on top of previously trained linear attention + sliding window models (e.g., LoLCATs) to significantly improve associative recall. As a result, LoLA extracts stronger language modeling capabilities from the same base model weights.
Enables Associative Recall. As a lightweight inference strategy, LoLA enables pass-key retrieval on up to 8K context lengths in needle-in-a-haystack tasks from the RULER benchmark [25]. With a 4.6x smaller cache than Llama-3.1 8B [20], our approach boosts accuracy from LoLCATs’ 0.6% to 97.4% at 4K context lengths with the same model weights. In our ablations, we show that this performance increase cannot be obtained from using a larger sliding window: global, sparse caching is essential.
Improves Language Modeling. LoLA improves the performance on zero-shot commonsense reasoning tasks from LM Evaluation Harness [17]. LoLA shows superior performance on many tasks among 1B and 8B parameter subquadratic architectures. This demonstrates that effective memory management can boost performance on language modeling tasks.
Reproducible on Consumer GPUs. LoLA achieves high performance on an academic budget. Nearly all of our method’s results can be reproduced on a single Nvidia RTX 4090.
@@ 2 Preliminaries
2.1 Softmax Attention
Transformers process a sequence of input tokens $\{\bm{x}_{t}\}_{t=1}^{n}$, for $\bm{x}_{t}\in\mathbb{R}^{d}$ [47]. The tokens are transformed into three distinct representations—queries, keys, and values—via trainable weight matrices $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}\in\mathbb{R}^{d\times d}$. For a given token $\bm{x}_{t}$, define
@@ 2.1 Softmax Attention
Causal attention uses the current query to recall past information from key-value pairs. The similarity between the query $\bm{q}_{t}$ and key $\bm{k}_{i}$ is denoted as $\alpha_{ti}\in(0,1)$. This similarity score determines how much value $\bm{v}_{i}$ is used for the current output token. The output token $\bm{y}_{t}$ is defined by
$\bm{y}_{t}=\sum_{i=1}^{t}\alpha_{ti}\,\bm{v}_{i}\in\mathbb{R}^{d},\quad\text{with}\quad\alpha_{ti}=\frac{\exp\left(\bm{q}_{t}^{\top}\bm{k}_{i}/\sqrt{d}\right)}{\sum_{j=1}^{t}\exp\left(\bm{q}_{t}^{\top}\bm{k}_{j}/\sqrt{d}\right)}.$ (2)
The collection of key-value pairs, $\{(\bm{k}_{i},\bm{v}_{i})\}_{i=1}^{t}$ , forms an associative memory bank [7]. Attention leverages this to learn a non-parametric key-to-value map at inference time. The output token is the predicted value associated with a given query. Softmax attention caches all key-value pairs to perform this non-parametric, or “look-up table”, operation. This leads to an unbounded KV-cache size of $\mathcal{O}(nd)$ with respect to sequence length $n$.
@@ 2.2 Linear Attention
Linear attention performs a low-rank approximation of the exponential dot product kernel [27]. This enables models to maintain constant-size memory footprints even for infinite sequence lengths. With this kernel approximation,
$\phi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{D},\quad\text{such that}\quad\exp\left(\frac{\bm{q}_{t}^{\top}\bm{k}_{j}}{\sqrt{d}}\right)\approx\phi(\bm{q}_{t})^{\top}\phi(\bm{k}_{j}),$ (3)
@@
$\bm{y}_{t}^{\top}=\sum_{i=1}^{t}\frac{\exp\left(\bm{q}_{t}^{\top}\bm{k}_{i}/\sqrt{d}\right)\;\bm{v}_{i}^{\top}}{\sum_{j=1}^{t}\exp\left(\bm{q}_{t}^{\top}\bm{k}_{j}/\sqrt{d}\right)}\approx\sum_{i=1}^{t}\frac{\phi(\bm{q}_{t})^{\top}\phi(\bm{k}_{i})\;\bm{v}_{i}^{\top}}{\sum_{j=1}^{t}\phi(\bm{q}_{t})^{\top}\phi(\bm{k}_{j})}$ $=\frac{\phi(\bm{q}_{t})^{\top}\left(\sum_{j=1}^{t}\phi(\bm{k}_{j})\;\bm{v}_{j}^{\top}\right)}{\phi(\bm{q}_{t})^{\top}\left(\sum_{j=1}^{t}\phi(\bm{k}_{j})\right)}$
@@
This creates a hidden state matrix $\mathbf{H}_{t}\in\mathbb{R}^{D\times d}$ as the sum of key-value outer products. This effectively bounds the memory cost to $\mathcal{O}(Dd)$, constant with respect to sequence length $n$. The hidden dimension size $D$ controls the approximation quality at the cost of computational efficiency. As a linear RNN, the hidden state $\mathbf{H}_{t}$ and normalization state $\bm{s}_{t}\in\mathbb{R}^{D}$ can be computed recurrently,
@@
In this formulation, linear attention treats each associative memory as a rank-one outer product. Rather than building a look-up table, linear attention parameterizes the key-to-value map as a linear function. This is viewed as an online learning problem, where the hidden state fits to the past KV-pairs in context [50].
@@ 2.3 Efficient Training of Linear Attention
To reduce the training costs, LoLCATs [56] and others [6, 49, 5] recycle large pretrained transformers into linear attention models with knowledge distillation [24]. These approaches minimize the difference between the pretrained transformer’s output $\bm{y}$ (i.e., the teacher) and linear attention’s output $\widehat{\bm{y}}$ (i.e., the student). In particular, LoLCATs uses a trainable nonlinear map for $\phi:\mathbb{R}^{d}\to\mathbb{R}^{D}$, constructed as
@@
with learnable weights $\bm{w}_{i}\in\mathbb{R}^{d}$ [57]. This distillation approach freezes all other parameters, adjusting $\phi$ to minimize the loss
@@
After attention distillation, the whole model is finetuned with LoRA [26]. Overall, this procedure only requires 40 million training tokens from the Alpaca dataset [46], grouped in 1024-long sequences.
@@ 2.4 Disadvantages of Previous Approaches
Even with distillation, linear attention models struggle to accurately mimic the behavior of softmax attention. Initial work in linear attention proposed nonlinear query and key activations to improve the exponential dot product approximation [11, 57]. These methods fall short as they are essentially performing a low-rank approximation of the infinite-rank exponential dot product kernel. In Appendix C, we show that the exponential dot product kernel has slowly decaying singular values for simple data distributions. This implies that high-dimensional hidden states may be required for modest approximation errors.
Recent approaches attempt to address the poor performance of the low-rank approximation in linear attention by augmenting it with sliding window attention [1, 56]. These methods compute a finite number of recent tokens in a window with softmax attention and compute the rest with linear attention. Since natural language contains a significant amount of local information, this hybrid approach nearly recovers the performance of pretrained transformers on short-context tasks. For longer sequences, however, these methods struggle to recall important information that falls outside the window and in the hidden state. We show in Table 1 that these models cannot perform associative recall on simple needle-in-a-haystack tasks. Other forms of sparse attention may be required [10].
@@ 3 Mitigating Memory Collisions with Sparse Caching
Associative memory systems with perfect recall allow stored keys to retrieve their own value. Similar to (2.2), we use a given key as a query in linear attention’s forward pass to verify this behavior with
$\frac{\phi(\bm{k}_{i})^{\top}\mathbf{H}_{t}}{\phi(\bm{k}_{i})^{\top}\bm{s}_{t}}=\frac{\phi(\bm{k}_{i})^{\top}\sum_{j=1}^{t}\phi(\bm{k}_{j})\,\bm{v}_{j}^{\top}}{\phi(\bm{k}_{i})^{\top}\sum_{j=1}^{t}\phi(\bm{k}_{j})}=\bm{v}_{i}^{\top}.$ (8)
In practice, we observe that "memory collisions" often occur. This means non-orthogonal keys interfere with each other when forming the hidden state $\mathbf{H}_{t}$. We define the self-recall score of a given key-value pair as
$\text{Score}(\bm{k},\bm{v}\,|\,\mathbf{H}_{t},\bm{s}_{t})=\left\|\frac{\phi(\bm{k})^{\top}\mathbf{H}_{t}}{\phi(\bm{k})^{\top}\bm{s}_{t}}-\bm{v}\right\|_{2}.$ (9)
This determines the error between the predicted value for a given key and the ground truth value. In this work, we demonstrate that a “self-recall check” is useful to see if a key-value pair is properly stored. LoLA uses this scoring approach to decide which KV pairs should be stored separately in full-rank. Large scores indicate the severity of the memory collision. As a result, LoLA maintains the top-scoring KV pairs in a sparse cache. This limits the corruption of past memories and improves associative recall.
@@
LoLA addresses the limitations of previous linear attention mechanisms by integrating a sparse caching strategy at inference time. This method employs three memory systems to approximate the softmax kernel effectively and better manage the hidden state
@@
    Linear Attention utilizes a finite-rank approximation to store an infinite amount of generic tokens.
@@
    Sliding Window Attention provides full-rank attention scores for local context.
@@
    Sparse Caching identifies and stores key-value pairs that are challenging to remember, maintaining high-fidelity representations for long-term dependencies.
Figure 1: LoLA stores past KV pairs in three forms memory.
LoLA mitigates memory collisions by determining which key-value pairs disagree with the current hidden state’s associative map. At every iteration, LoLA scores the KV pair leaving the sliding window and re-scores all pairs in the sparse cache. Re-scoring is needed since the self-recall error is dependent on the current hidden state. A KV pair could become “easier to remember” in the future, after hidden state updates.
We define the pairs that are scored at time $t$ as
@@
where $G_{t-1}$ is the set of KV pairs in the sparse cache at time $t-1$. Also, $\eta$ is the maximum number of pairs in the sliding window. Then, we update the sparse cache by taking the top-$\lambda$ scoring pairs in $\mathcal{E}_{t}$, i.e,
$G_{t}=\operatorname*{arg\,max}_{G\subset\mathcal{E}_{t}:|G|=\lambda}\sum_{(\bm{k},\bm{v})\in G}\text{Score}(\bm{k},\bm{v}\,|\,\mathbf{H}_{t},\bm{s}_{t}).$ (11)
@@
When the input sequence is available ahead of time (e.g., prefill), LoLA is accelerated with parallelization [52]. By partitioning the input sequence into chunks of size $C$, we can compute intra-chunk operations in parallel with dense matmuls. This reduces the number of recurrent iterations by a factor of $C$ while preserving the constant-memory cost that motivates LoLA.
LoLA computes softmax attention with the current chunk of queries and past two chunks of KV-pairs. For small chunk sizes, softmax attention is almost equally efficient to linear attention. Artificially limiting softmax within a chunk will not improve efficiency, only hurt performance. The past two chunks of KV-pairs are concatenated with the sparse cache in order to compute a single FlashAttention [13] pass per chunk of queries. For the linear attention outputs, all queries within the chunk share the same hidden state.
After computing the past chunk of output tokens, we evict the oldest KV-pairs in the window, sending them to the hidden state or sparse cache. All of the evicted and sparse cache pairs are scored with the self-recall error (9). The $\lambda$ pairs with the largest errors in the eligible set,
@@
will remain in the sparse cache. The remainder are integrated into the hidden state through the standard outer product update (5).
@@
The given hardware setup dictates the total fixed cache size of LoLA, but the ratio of chunk size to sparse cache size depends on the application. Increasing the chunk size $C$ reduces the number of recurrent iterations and overhead costs from sparse caching. This puts more emphasis on local KV-pairs since there will be a larger portion of sliding window attention. Increasing the sparse cache size $\lambda$ will better mitigate collisions in the hidden state and improve long context recall. For simplicity in our recall experiments, we choose $C=\lambda$ and report various total cache sizes.

### v2

@@ Abstract
[…]
@@ 1 Introduction
Transformer-based large language models (LLMs) rely on storing all past tokens in an ever-growing key-value (KV) cache (Vaswani et al., 2017). This allows future query tokens to access past memories with associative recall, which enables in-context learning (Olsson et al., 2022). Since no previous information is discarded, the KV cache continues to grow with context length. This eventually leads to a memory bottleneck on long context tasks, such as lifelong in-context learning. As a result, transformers cannot condition next token predictions on arbitrarily long sequences.
Alternative architectures to transformers have been proposed—such as Mamba (Gu & Dao, 2024), DeltaNet (Schlag et al., 2021), linear attention (Katharopoulos et al., 2020), and others (Yang et al., 2024a; Behrouz et al., 2024; Sun et al., 2024)—to reduce the compute complexity from quadratic to linear. Additionally, these approaches reduce the memory cost from linear to constant. In particular, linear attention removes the exponential dot product in softmax (Katharopoulos et al., 2020). This effectively collapses the unbounded KV-cache into a fixed-size matrix, which corresponds to a recurrently formed hidden state (i.e., a linear RNN). This constructs a linear associative memory map from keys to values. Past memories can be recalled through a vector-matrix product of an incoming query vector and the hidden state matrix. Linear attention enables constant-cost prediction per token when conditioned on arbitrarily long contexts.
While efficient and flexible, linear attention architectures lag behind transformers in terms of memory capacity. This is largely noticeable on tasks leveraging in-context learning (Paperno et al., 2016; Hsieh et al., 2024). The removal of the exponential dot product allows for non-orthogonal keys to interfere with the hidden state’s learned key-to-value map. This interference—denoted as a memory collision—impairs associative recall. Previous work used nonlinear query and key activations to improve the exponential dot product approximation (Choromanski et al., 2020; Zhang et al., 2024). However, these attempts are essentially performing a low-rank approximation of the infinite-rank exponential dot product kernel.
Additional use of sparse attention (Chen et al., 2021) can improve linear attention’s recall; however, current hybrid approaches only focus on local information with sliding window attention (Arora et al., 2024; Zhang et al., 2025a; Lan et al., 2025; Van Nguyen et al., 2025). These approaches can struggle to recall critical, long-term facts that fall outside the window.
@@
How can long term associative memory for subquadratic language models be improved?
@@
We present LoLA: Low-rank Linear Attention with sparse caching. LoLA is a novel, training-free inference strategy that boosts the performance of hybrid linear attention layers. LoLA distributes historical tokens into three forms of memory: (i) recent KV pairs are stored in a sliding window cache, (ii) difficult-to-memorize pairs in a sparse global cache, and (iii) all other pairs are placed in a recurrent hidden-state matrix via linear attention. LoLA performs a self-recall check to see which KV pairs disagree with the current hidden state’s linear associative map. LoLA sparsely caches the interfering memories in full rank. The selection mechanism effectively mitigates memory collisions with a small, constant-sized cache. This inference strategy can be applied on top of previously trained linear attention + sliding window models (e.g., LoLCATs) to significantly improve associative recall. As a result, LoLA extracts stronger language modeling capabilities from the same base model weights.
Utilizes Self-Recall Error. We introduce an importance metric for key-value pairs to reduce memory collisions in the hidden state of linear attention. This is computed by determining if a key can recall its own value with linear attention. In our ablations, we show that this performance increase cannot be obtained from using a larger sliding window or other sparse attention metrics: self-recall is essential.
Enables Associative Recall. As a lightweight inference strategy, LoLA enables pass-key retrieval on up to 8K context lengths in needle-in-a-haystack tasks from the RULER benchmark (Hsieh et al., 2024). With a 4.6x smaller cache than Llama-3.1 8B (Grattafiori et al., 2024), our approach boosts accuracy from LoLCATs’ 0.6% to 97.4% at 4K context lengths with the same model weights.
Improves Language Modeling. LoLA shows superior performance on zero-shot commonsense reasoning tasks among 1B and 8B parameter subquadratic architectures. This demonstrates that effective memory management can boost language modeling performance.
@@ 2 Preliminaries
2.1 Softmax Attention as a Nonparametric, Online Learner
Transformers process a sequence of input tokens $\{\bm{x}_{t}\}_{t=1}^{n}$, for $\bm{x}_{t}\in\mathbb{R}^{d}$ (Vaswani et al., 2017). For each attention head, the input tokens are transformed into three distinct representations—queries, keys, and values—via trainable weight matrices $\mathbf{W}_{q},\mathbf{W}_{k}\in\mathbb{R}^{d\times d_{k}}$ and $\mathbf{W}_{v}\in\mathbb{R}^{d\times d_{v}}$. For a given token $\bm{x}_{t}$, define
@@ 2.1 Softmax Attention as a Nonparametric, Online Learner
Causal attention uses the current query to recall past information from key-value pairs. The similarity between the query $\bm{q}_{t}$ and key $\bm{k}_{i}$ is denoted as $\alpha_{ti}\in(0,1)$. This similarity score determines how much value $\bm{v}_{i}$ is used for the current output token at time $t$. The output token $\bm{y}_{t}$ is defined by
$\bm{y}_{t}=\sum_{i=1}^{t}\alpha_{ti}\,\bm{v}_{i}\in\mathbb{R}^{d_{v}},\quad\text{with}\quad\alpha_{ti}=\frac{\exp\left(\bm{q}_{t}^{\top}\bm{k}_{i}/\sqrt{d_{k}}\right)}{\sum_{j=1}^{t}\exp\left(\bm{q}_{t}^{\top}\bm{k}_{j}/\sqrt{d_{k}}\right)}.$ (2)
We view softmax attention as a nonparametric function $m_{t}:\mathbb{R}^{d_{k}}\to\mathbb{R}^{d_{v}}$ that fits to the past context at inference time. This online function $m_{t}$ learns to map keys to their associated value in context with $m_{t}(\bm{k}_{i})\approx\bm{v}_{i}$ for $(\bm{k}_{i},\bm{v}_{i})\in\{(\bm{k}_{i},\bm{v}_{i})\}_{i=1}^{t}$. Then, $m_{t}$ applies the learned transformation to the query, $\bm{y}_{t}=m_{t}(\bm{q}_{t})$. The set of past key-value pairs forms an online "training set" of input-output labels. The query acts as an unsupervised "test set".
Softmax attention caches all past key-value pairs to perform this non-parametric, or “look-up table”, operation. Since the function complexity scales with the context length, softmax attention can flexibly learn new context without forgetting past key-value associations. However, this process leads to an unbounded KV-cache size that scales linearly with sequence length $n$. Ultimately, this operation cannot be used for extremely long context scenarios, such as lifelong learning.
@@ 2.2 Linear Attention
To bound inference costs, linear attention methods replace the exponential dot product kernel (Katharopoulos et al., 2020) with a low-rank approximation. This enables models to maintain constant-size memory footprints even for infinite sequence lengths. With this replacement,
$\exp\left(\frac{\bm{q}_{t}^{\top}\bm{k}_{j}}{\sqrt{d_{k}}}\right)\approx\phi(\bm{q}_{t})^{\top}\phi(\bm{k}_{j}),\quad\text{for}\quad\phi:\mathbb{R}^{d_{k}}\rightarrow\mathbb{R}^{D},$ (3)
@@
$\bm{y}_{t}^{\top}=\sum_{i=1}^{t}\frac{\exp\left(\bm{q}_{t}^{\top}\bm{k}_{i}/\sqrt{d_{k}}\right)\;\bm{v}_{i}^{\top}}{\sum_{j=1}^{t}\exp\left(\bm{q}_{t}^{\top}\bm{k}_{j}/\sqrt{d_{k}}\right)}\approx\sum_{i=1}^{t}\frac{\phi(\bm{q}_{t})^{\top}\phi(\bm{k}_{i})\;\bm{v}_{i}^{\top}}{\sum_{j=1}^{t}\phi(\bm{q}_{t})^{\top}\phi(\bm{k}_{j})}$ $=\frac{\phi(\bm{q}_{t})^{\top}\left(\sum_{j=1}^{t}\phi(\bm{k}_{j})\;\bm{v}_{j}^{\top}\right)}{\phi(\bm{q}_{t})^{\top}\left(\sum_{j=1}^{t}\phi(\bm{k}_{j})\right)}$
@@
This creates a hidden state matrix $\mathbf{H}_{t}\in\mathbb{R}^{D\times d_{v}}$ as the sum of key-value outer products. This effectively bounds the memory cost to $\mathcal{O}(Dd_{v})$, constant with respect to sequence length $n$. The hidden dimension size $D$ controls the approximation quality at the cost of computational efficiency. As a linear RNN, the hidden state $\mathbf{H}_{t}$ and normalization state $\bm{s}_{t}\in\mathbb{R}^{D}$ can be computed recurrently,
@@
In this formulation, linear attention stores each observation, or KV-pair, as a rank-one outer product. Rather than building a look-up table, linear attention parameterizes the key-to-value map as a linear function. While this approach is efficient, linear attention falls short in memory capacity as the number of orthogonal key-value pairs is bounded by the rank of $\mathbf{H}_{t}$. “Memory collisions” (Yang et al., 2024a) occur when new hidden state updates overwrite past key-value associations. This prevents the linear map from accurately modeling the context.
@@ 2.3 Efficient Training of Linear Attention
To reduce training costs, LoLCATs (Zhang et al., 2025a) and others (Bick et al., 2024; Wang et al., 2024; Bick et al., 2025; Goldstein et al., 2025) recycle large pretrained transformers into linear attention models with knowledge distillation (Hinton et al., 2015). These approaches minimize the difference between the pretrained transformer’s output $\bm{y}$ (i.e., the teacher) and linear attention’s output $\widehat{\bm{y}}$ (i.e., the student). In particular, LoLCATs uses a trainable nonlinear map for $\phi:\mathbb{R}^{d_{k}}\to\mathbb{R}^{D}$, constructed as
@@
with learnable weights $\bm{w}_{i}\in\mathbb{R}^{d_{k}}$ (Zhang et al., 2024). This distillation approach freezes all other parameters, adjusting $\phi$ to minimize the loss
@@
After attention distillation, the whole model is finetuned with LoRA (Hu et al., 2022). Overall, this procedure only requires 40 million training tokens from the Alpaca dataset (Taori et al., 2023), grouped in 1024-long sequences.
@@ 2.4 Disadvantages of Previous Approaches
Even with distillation, linear attention models struggle to accurately mimic the behavior of softmax attention. Initial work in linear attention proposed nonlinear query and key activations to improve the exponential dot product approximation (Choromanski et al., 2020; Zhang et al., 2024). These methods fall short as they are essentially performing a low-rank approximation of the infinite-rank exponential dot product kernel. In Appendix E, we show that the exponential dot product kernel has slowly decaying singular values for simple data distributions. This implies that high-dimensional hidden states may be required for modest approximation errors.
Recent approaches attempt to address the poor performance of the low-rank approximation in linear attention by augmenting it with sliding window attention (Arora et al., 2024; Zhang et al., 2025a). These methods compute a finite number of recent tokens in a window with softmax attention and compute the rest with linear attention. Since natural language contains a significant amount of local information, this hybrid approach nearly recovers the performance of pretrained transformers on short-context tasks. For longer sequences, however, these methods struggle to recall important information that falls outside the window and in the hidden state. We show in Table 1 that these models cannot perform associative recall on simple needle-in-a-haystack tasks. Other forms of sparse attention may be required alongside "low-rank" attention (Chen et al., 2021).
@@ 3 Mitigating Memory Collisions with Sparse Caching
Identifying Difficult-to-Remember KV Pairs.
As our base assumption, strong associative memory systems should allow keys to retrieve their associated values. Online functions $m_{t}$ with perfect recall interpolate the online training set, defined as
$m_{t}(\bm{k}_{i})=\bm{v}_{i},\quad\forall i\leq t$ (8)
For perfect recall in linear attention, equation 8 translates to
$m_{t}(\bm{k}_{i})^{\top}=\frac{\phi(\bm{k}_{i})^{\top}\mathbf{H}_{t}}{\phi(\bm{k}_{i})^{\top}\bm{s}_{t}}=\frac{\phi(\bm{k}_{i})^{\top}\sum_{j=1}^{t}\phi(\bm{k}_{j})\,\bm{v}_{j}^{\top}}{\phi(\bm{k}_{i})^{\top}\sum_{j=1}^{t}\phi(\bm{k}_{j})}=\bm{v}_{i}^{\top}.$ (9)
In practice, however, "memory collisions" prohibit equation 9 from holding. Non-orthogonal keys interfere with each other when forming the hidden state $\mathbf{H}_{t}$. We measure the Self-Recall Error (SRE) to measure how well a past key can retrieve its associated value with
$\text{SRE}(\bm{k},\bm{v}\,|\,\mathbf{H}_{t},\bm{s}_{t})=\left\|\frac{\phi(\bm{k})^{\top}\mathbf{H}_{t}}{\phi(\bm{k})^{\top}\bm{s}_{t}}-\bm{v}\right\|_{2}=\left\|\hat{\bm{v}}-\bm{v}\right\|_{2}.$ (10)
This determines the error between the predicted value $\hat{\bm{v}}$ for a given key $\bm{k}$ and the ground truth value $\bm{v}$.
@@
We propose LoLA: a training-free inference strategy that boosts the performance of hybrid linear attention models. LoLA addresses the limitations of previous linear attention mechanisms by integrating a sparse caching strategy at inference time. This method employs three memory systems to store long term associations
@@
    Linear Attention utilizes a finite-rank approximation to store an infinite amount of tokens.
@@
    Sliding Window Attention provides full-rank attention scores for finite, local context.
@@
    Sparse Caching identifies and stores key-value pairs that are challenging to remember, preventing memory collisions in linear attention.
Figure 1: LoLA stores past KV pairs in three forms memory for each attention head.
LoLA uses the self-recall error, equation 10, to decide which KV pairs should be stored separately in full-rank. Large errors indicate the severity of the memory collision. As a result, LoLA keeps the KV pairs with the largest error in a sparse cache. This limits the corruption of past memories and improves associative recall.
Since only a finite amount of tokens can be stored at each time step to maintain efficiency, LoLA performs a greedy scoring approach. At every iteration, LoLA scores the KV pairs leaving the sliding window and re-scores all pairs currently stored in the sparse cache. At each timestep, the pairs with the lowest error are moved to the linear hidden state indefinitely. Re-scoring pairs in the sparse cache is essential since the SRE is dependent on the current hidden state. For example, a KV pair could become more aligned with the hidden state in the future after a few updates. In the generation implementation of LoLA, we define the set of pairs that are scored at time $t$ as
@@
where $G_{t-1}$ is the set of KV pairs in the sparse cache at time t1t-1. Here, $\eta$ is the maximum number of pairs in the sliding window. We update the sparse cache by selecting the top-$\lambda$ errors in $\mathcal{E}_{t}$, i.e,
$G_{t}=\operatorname*{arg\,max}_{G\subset\mathcal{E}_{t}:|G|=\lambda}\sum_{(\bm{k},\bm{v})\in G}\text{SRE}(\bm{k},\bm{v}\,|\,\mathbf{H}_{t},\bm{s}_{t}).$ (12)
@@
When the input sequence is available ahead of time (e.g., prefill), LoLA is accelerated with parallelization. By partitioning the input sequence into chunks of size $C$, we can compute intra-chunk operations in parallel with dense matmuls (Yang et al., 2024b). This reduces the number of recurrent iterations by a factor of $C$ while preserving the constant-memory cost that motivates LoLA.
LoLA computes softmax attention with the current chunk of queries and previous two chunks of KV-pairs. For small chunk sizes, softmax attention is almost equally efficient to linear attention. Artificially limiting softmax within a chunk will not improve efficiency, only hurt performance. The past two chunks of KV-pairs are concatenated with the sparse cache in order to compute a single FlashAttention (Dao et al., 2022) pass per chunk of queries. For the linear attention portion of the forward pass, all queries within the chunk share the same hidden state.
After computing the past chunk of output tokens, we evict the oldest chunk of KV-pairs in the window, sending them to the hidden state or sparse cache. All of the evicted and sparse cache pairs are scored with the SRE, equation 10. The $\lambda$ pairs with the largest errors in the eligible set,
@@
will remain in the sparse cache. The remainder are integrated into the hidden state through the standard outer product update as in equation 5.
@@
The given hardware setup dictates the total fixed cache size of LoLA, but the ratio of chunk size to sparse cache size depends on the application. Increasing the chunk or sliding window size reduces the number of recurrent iterations and overhead costs from sparse caching; however, this requires more VRAM. Increasing the sparse cache size $\lambda$ will better mitigate collisions in the hidden state and improve long context recall. In Appendix B, we explore this trade-off for various cache hyperparameters in an efficiency analysis. Specifically we measure the total VRAM use, Time-to-First-Token, and long context performance. Furthermore, we illustrate the bounded nature of LoLA, compared to vanilla transformers.